OpenStudy (anonymous):

For the functions f(x) = 5x − 3 and g(x) = x + 4, which composition produces the greatest output? f(g(x)) produces the greatest output. g(f(x)) produces the greatest output.<-----my ANSWER Both compositions produce the same output. Neither composition produces an output.

OpenStudy (anonymous):

i need help plzz

OpenStudy (australopithecus):

These are linear equations, remember y = mx + b g(f(x)) = (5x − 3) + 4 = 5x +1 f((g(x)) = 5(x + 4) − 3 = 5x + 17

OpenStudy (australopithecus):

They both have the same slope

OpenStudy (australopithecus):

the only thing that differs between them is the y-intercept

OpenStudy (anonymous):

so its C

OpenStudy (anonymous):

WAIT NVM. its A

OpenStudy (australopithecus):

yup http://www.wolframalpha.com/input/?i=y+%3D+5x+%2B1%2C+y+%3D+5x+%2B+17+ here just to confirm your answer is correct to you look at their graphs

OpenStudy (anonymous):

oo ok. can u help me with one more problem

OpenStudy (anonymous):

Which of the following expressions is the inverse of the function y equals quantity x minus 2 divided by 3? y = 3x + 2 y = 2x + 3 y equals quantity x minus 3 divided by 2 y equals quantity x plus 2 divided by 3

OpenStudy (australopithecus):

First write out the equation, then to find the inverse set y to x and x to y then solve for y

OpenStudy (anonymous):

x-2/3

OpenStudy (australopithecus):

well, y = x - 2/3

OpenStudy (australopithecus):

now replace y with x and x with y and solve for y and you have the inverse of that function

OpenStudy (anonymous):

im confused

OpenStudy (anonymous):

wait so its B

OpenStudy (australopithecus):

If you had y = x + 2 to find the inverse x = y + 2 solve for y y = x - 2

OpenStudy (anonymous):

Y=3X+2

OpenStudy (anonymous):

THXXX ALOT